diferenciate with respect to x and y U=25X^2/5Y^3/5  


1
Expert's answer
2022-02-03T12:17:00-0500

"\frac {\delta U} { \delta X}=25\times\frac{2}{5}(X^\frac{-3}{5}Y^\frac{3}{5 })"

"\frac {\delta U} { \delta X}=10(X^\frac{-3}{5}Y^\frac{3}{5 })"


"\frac {\delta U} { \delta Y}=25\times\frac{3}{5}(X^\frac{2}{5}Y^\frac{-2}{5 })"


"\frac {\delta U} { \delta Y}=15(X^\frac{3}{5}Y^\frac{-2}{5 })"